OpenStudy (anonymous):

Simplify the expression -2 + 3(2x – 1) + 6x

OpenStudy (anonymous):

8x 12x + 5 12x - 5 8x - 3

OpenStudy (anonymous):

multiply 3 to 2x+1

OpenStudy (anonymous):

we will get -2 + 6x - 3 + 6x , u may do last step ...

OpenStudy (anonymous):

12x -5 ( correct answer )

OpenStudy (anonymous):

Expand the brackets first: =3 * 2x -3*1 =6x-3 Simplify: =-2 +6x-3+6x = -5 +12 x Rearrange so that variable is first: =12x -5